Please help me solve: A woman has $1.70 in dimes and nickels. She has 2 more dimes than nickels. How many nickels does she have?
Answer by mananth(16949)

A woman has $1.70 in dimes and nickels. She has 2 more dimes than nickels. How many nickels does she have?
let the number of nickels be x
then the number of dimes she will have is x+2
The value of nickels will be 5x
and the value of dimes will be 10(x+2)
The sum of both the values is = 170 cents
5x+10(x+2) = 170
5x+10x+20 = 170
15x=170-20
15x= 150
x= 150/15 = 10
She has 10 nickels
and 12 dimes ( x+2)